Differentiate the following without solving for y in terms of x.

  1. x2+10x+y2+17=0
  2. x2+y2-2x+4y=3
  3. x2 +y2+2y=45
  4. x2-4x+y2+2y=15
  5. x2-2x+y2+2y=23
1
Expert's answer
2022-04-12T05:30:24-0400
  1. "2x+10+2y*y'=0"
  2. "2x+2y*y'-2+4y'=0"
  3. "2x+2y*y'+2y'=0"
  4. "2x-4+2y*y'+2y'=0"
  5. "2x-2+2y*y'+2y'=0"